OpenStudy (alexh107):

Rewrite with only sin x and cos x. sin 2x - cos x

OpenStudy (alexh107):

Is this just sin x because 2x-x =x?

OpenStudy (mrnood):

no - you cannot do that look up in your text the identity for sin(2x) you will find an equation which gives you sin 2terms of sinx and cosx Then substitute that in the above equation and simplify it

OpenStudy (mrnood):

no - you cannot do that look up in your text the identity for sin(2x) you will find an equation which gives you sin 2x in terms of sinx and cosx Then substitute that in the above equation and simplify it

OpenStudy (alexh107):

sin 2x + 2 sin x cos?

OpenStudy (alexh107):

sin 2x = 2 sin x cos x*

OpenStudy (mrnood):

yes that is the correct substitution. So put that in the original and then simplify (by factoring out cos x)

OpenStudy (alexh107):

\[2 \sin x \cos x - \cos x = 2 \sin x\]?

OpenStudy (alexh107):

@MrNood

OpenStudy (mrnood):

no that's not th ecorrect simplification you should 'factor out' cos x cosx( ???? )

OpenStudy (alexh107):

cos x (2 sin x - 1)

OpenStudy (alexh107):

@MrNood

OpenStudy (mrnood):

correct

OpenStudy (alexh107):

Thank you!